# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ IV-D MLLM-Assisted UAV Navigation and Collaboration
At a broader level, Liu et al. [76] introduce AerialVLN, a benchmark task that redefines UAV navigation in open, continuous 3D urban spaces. Unlike ground-based VLN, AerialVLN requires UAVs to navigate freely using first-person visual input while handling altitude variation and ambiguous spatial references. This task formalizes key challenges in cross-modal grounding and 3D spatial reasoning that contemporary frameworks such as UAV-VLN, SkyVLN, and AeroDuo aim to address. A comparative summary is provided in Table XI.
Overall, recent VLN advances demonstrate that tightly integrating LLM-assisted reasoning with visual perception, control-aware execution, and multi-UAV collaboration enables UAVs to translate natural-language instructions into real-time, context-aware flight behavior. These developments establish a strong foundation for future research in scalable, language-driven aerial navigation.
Table XI: Comparison of VLN frameworks for UAVs, highlighting innovations in end-to-end reasoning, safety-aware control, instruction robustness, multi-UAV collaboration, and benchmark development for continuous 3D aerial navigation.
@@
[76] Proposes a new VLN benchmark for UAVs in continuous 3D outdoor urban spaces. Task definition for continuous, open 3D navigation with first-person visual perceptions. Introduces challenges in 3D spatial reasoning, language understanding, and cross-modal alignment. Addresses the gap in aerial agent navigation tasks and focuses on complex urban environments.
@@
IV-E LLM/VLM-Assisted High-Level Mission Planning, Task Reasoning, and Operational Intelligence
@@
Recent research has increasingly focused on enabling UAVs to autonomously interpret visual scenes, understand natural language, and perform time-critical operations. Yaqoot et al. [128] exemplify this trend with UAV-VLRR, a rapid SAR framework that integrates a multimodal vision–language reasoning system using ChatGPT-4o and Molmo with a high-performance NMPC. By directly translating language commands and aerial imagery into actionable target coordinates and executing point-to-point NMPC control with integrated obstacle avoidance, UAV-VLRR eliminates the need for preplanned trajectories. Experimental results show significant speed improvements over both human pilots and conventional autopilots, highlighting the effectiveness of combining MLLMs with real-time control for agile, safety-critical missions.
Complementing rapid execution at the control level, Sautenkov et al. [101] extend multimodal reasoning to strategic mission planning with UAV-VLA. Instead of directly controlling low-level flight, their system enables users to specify large-scale aerial missions using natural language, which an LLM interprets as high-level objectives. A vision–language model then identifies relevant targets in satellite imagery, and a final LLM module generates a complete flight plan. Although the resulting trajectories are slightly longer than those designed by expert operators, they are produced more than six times faster, demonstrating the potential of MLLMs to reduce operator workload and streamline global-scale mission design significantly. Beyond navigation and planning, multimodal AI has also demonstrated strong potential for real-time incident detection and situational awareness. Eesaar et al. [24] present an autonomous UAV system that integrates a DJI Matrice 30T with an end-to-end multimodal pipeline for emergency response. In this system, YOLOv12n detects accidents and fires in real-time aerial video, LLaVA-OneVision-Qwen2 generates detailed scene-level descriptions, and GPT-4o mini produces concise natural-language summaries for first responders. Field validation shows that this pipeline can rapidly transform raw aerial imagery into actionable alerts, significantly improving response times and situational awareness in transportation and emergency management scenarios.
@@
IV-F LLM/VLM-Driven Autonomous Control and Real-Time Flight Decision-Making
LLMs and VLMs in UAV systems pose a critical challenge: enabling autonomous agents to accurately infer human intent, interpret complex, dynamic environments, and make safe, context-aware decisions despite the uncertainty and potential hallucinations inherent in modern foundation models. Recent advances address these challenges by combining structured prompt execution mechanisms with lightweight, task-specific vision models, improving reliability and reducing erroneous reasoning. Multimodal architectures further enhance human–machine collaboration by generating navigation commands with natural-language explanations and confidence estimates, enabling operators to supervise and guide UAV behavior in real time. Meanwhile, efficient onboard autonomy frameworks leverage compact, fine-tuned LLMs and advanced perception modules for object detection and depth estimation, enabling robust autonomous decision-making even under stringent computational constraints.
@@ IV-F LLM/VLM-Driven Autonomous Control and Real-Time Flight Decision-Making
Building on the theme of reliability while emphasizing human collaboration, Krupavs et al. [55] introduce a multimodal architecture for human–machine collaborative UAV navigation. Their system uses GPT-4.1-nano for multimodal reasoning and generates navigation commands along with natural-language explanations and confidence scores, which are presented to operators through a touch-based approval interface. This design enables a closed-loop interaction paradigm in which operators can accept, reject, or modify model-generated actions. Experimental results show high levels of user trust and explainability, with approximately 90% approval of AI-generated instructions. However, observed response latency and imperfect calibration between confidence scores and actual performance highlight ongoing challenges in deploying explainable VLMs for real-time aerial control.
@@
Extending these ideas to goal-directed navigation, Zhang et al. [135] propose VLFly, a unified vision–language navigation framework explicitly designed for UAVs that avoids the common reliance on discrete action spaces and external localization systems. Instead, VLFly directly outputs continuous velocity commands from monocular camera input, integrating an LLM-assisted instruction encoder, a VLM-based goal retriever, and a trajectory planner for real-time waypoint generation. Evaluated across diverse simulation environments without task-specific fine-tuning, VLFly consistently outperforms previous approaches. Real-world flight experiments in both indoor and outdoor settings further confirm its robust open-vocabulary grounding and its ability to interpret abstract or indirect language instructions, highlighting its suitability for zero-shot deployment in unconstrained UAV navigation scenarios. Table XII summarizes challenges and solution directions in LLM/VLM-Based UAV Autonomy.
@@
Table XII: Key challenges in LLM/VLM-based UAV autonomy—including model uncertainty, interpretability, resource constraints, environmental perception, and intent understanding—paired with solution approaches, required capabilities, and desired operational outcomes.
@@
IV-G Semantic Understanding, Analysis, and Interpretation
As UAVs increasingly operate in complex environments and regulated airspaces, traditional perception pipelines struggle to deliver the semantic understanding and real-time reasoning required for robust autonomy. Emerging multimodal foundation models offer a promising alternative, enabling UAVs to perform advanced inference, threat assessment, and scene understanding across diverse tasks. Applications include low-altitude airspace security, where data from multiple sensor modalities inform defensive decision-making, and precision agriculture, where vision–language models convert aerial imagery into actionable crop intelligence. These advances highlight the growing versatility of multimodal AI in UAV systems, supporting adaptive, efficient, and generalizable environmental intelligence. Fig. 9 visualizes LLM/VLM-assisted semantic Understanding, Analysis, and Interpretation.
As multimodal large models gain traction in UAV-related domains, recent research has begun to explore their potential beyond navigation and control, extending into higher-level reasoning for security and remote sensing applications. Lei et al. [57] exemplify this shift in low-altitude airspace security by proposing a system that integrates MLLMs with heterogeneous perception data to infer the intent of non-cooperative UAVs. By fusing inputs from RGB cameras, infrared sensors, and radar, the system constructs structured representations, which are processed by an MLLM that leverages contextual knowledge and generative reasoning to assess adversarial behavior and recommend defensive actions. A confrontation scenario demonstrates the feasibility of this approach. The authors also identify key challenges – including robust communication, scalability to multi-UAV scenarios, and effective edge–cloud collaboration – that outline future directions for intelligent and adaptive airspace defense architectures.
@@ IV-G Semantic Understanding, Analysis, and Interpretation
IV-H Key Findings and Insights
This section consolidates key observations and empirical findings on the deployment of Multimodal Large Language Models (MLLMs) in UAV applications. Emphasis is placed on quantitative evidence, benchmark results, and UAV-specific limitations that directly affect real-world reliability, efficiency, and safety as sequel.
IV-H1 Benchmarking MLLMs for Multi-UAV Collaborative Perception
AirCopBench [131] is the first benchmark explicitly designed to evaluate MLLMs in multi-UAV collaborative perception under degraded sensing conditions. The benchmark contains over 2,900 multi-view aerial images and more than 14,600 VQA samples, covering four task categories (Scene Understanding, Object Understanding, Perception Assessment, and Collaborative Decision) across 14 task types. Evaluation of 40 state-of-the-art MLLMs reveals a substantial performance gap: the best-performing model achieves only 59.17% accuracy, lagging human performance by more than 24% on average, with large variance across task categories. Fine-tuning results indicate that sim-to-real transfer can partially mitigate these gaps, highlighting the importance of UAV-specific data and training strategies for collaborative aerial perception.
IV-H2 Speculative Decoding for Efficient LLM Inference
Inference latency remains a critical bottleneck for UAV deployment, as conventional autoregressive decoding requires one forward pass per generated token. Speculative decoding improves efficiency by allowing multiple candidate tokens to be generated and verified in parallel within a single forward pass, significantly reducing end-to-end latency while preserving output correctness. Recent approaches such as EAGLE-3 further eliminate the need for a separate draft model by exploiting internal hidden states of the target model. Although primarily evaluated in general-purpose LLM settings, these techniques are highly relevant to UAV platforms, where real-time responsiveness and energy efficiency are tightly constrained.
IV-H3 Long-Context Modeling in Multimodal Systems
Long-context modeling enables multimodal systems to reason over extended inputs, but scaling context length introduces challenges including quadratic attention complexity, positional encoding degradation, and mid-context retrieval failure. Recent solutions—such as efficient attention mechanisms, memory-augmented architectures, and chunked processing—improve scalability, while data-centric strategies (e.g., synthetic long-context generation and short-to-long alignment) enhance robustness. However, for UAV applications involving long-duration missions or multi-stage reasoning, a gap persists between theoretical context length and effective usable context, limiting practical long-horizon multimodal reasoning [78].
IV-H4 Multi-Agent Framework for MLLMs Training
@@
IV-H5 Addressing Visual Hallucination in Multimodal LLMs
MLLMs have demonstrated remarkable capabilities across a wide range of tasks that integrate text and visual information. Despite these advances, a persistent challenge is Visual Hallucination (VH), the generation of outputs containing details inconsistent with the visual input. In other words, MLLMs may ”imagine” objects, attributes, or relationships that are absent from an image, particularly in tasks such as object recognition. This phenomenon undermines model reliability and limits their safe and effective deployment in UAV applications.
@@ IV-H5 Addressing Visual Hallucination in Multimodal LLMs
IV-H6 Balancing Efficiency and Reasoning Diversity in Vision-Language Models for UAV Applications
@@
Results in [38] demonstrate that UAV-oriented vision–language models can significantly outperform general-purpose models when adapted with domain-specific data and training strategies. The proposed UAV-VL-R1 model achieves strong reasoning performance while maintaining a compact footprint suitable for onboard deployment. Ablation studies reveal a key trade-off: while supervised fine-tuning (SFT) accelerates convergence and stabilizes early performance, excessive reliance on SFT reduces reasoning diversity, particularly in counting and numerical inference tasks. In contrast, the multi-stage GRPO framework enhances reasoning diversity and cross-task generalization through intra-group reward comparisons. The associated HRVQA-VL dataset, consisting of 50,019 samples organized into progressively challenging stages, further highlights that domain-specific data can outweigh sheer model scale, outperforming both 2B and 72B Qwen2-VL baselines under UAV constraints.
IV-H7 Understanding and Evaluating Hallucination Behavior in Aerial LVLMs
@@
Basak et al. [8] introduce AeroCaps, the first aerial-view image captioning dataset, comprising 1,256 human-annotated images. Their analysis shows that aerial perspectives significantly amplify hallucination rates, especially for object identity, spatial relations, and numerical descriptions, with error frequency increasing for longer captions. Notably, mitigation strategies such as object hints and LLM-based self-evaluation exhibit limited effectiveness, underscoring the difficulty of reliable hallucination detection. These findings indicate that current LVLMs remain insufficiently robust for real-world UAV deployment without further advances in dataset scale, evaluation rigor, and grounding-aware model design. The study emphasizes that, despite impressive performance on conventional benchmarks, current LVLMs are not yet sufficiently robust or trustworthy for deployment in real-world aerial applications, including automated surveillance, environmental monitoring, and disaster response. The results suggest that progress in this area will require not only improved model architectures and training strategies but also domain-specific datasets, rigorous evaluation protocols, and transparent reporting of failure modes. As aerial data becomes increasingly central to AI-driven decision-making, addressing these challenges will be critical to ensuring the safe deployment of AI.
In summary, based on Sections II and III, the multifaceted integration of LLMs and MLLMs into autonomous aerial systems necessitates an investigation of their ethical implications. LLMs are becoming central to UAV decision-making, perception, and control – capabilities with significant real-world consequences. Section II describes how LLMs, through prompt engineering and RAG, enable UAVs to perform complex tasks such as disaster response and navigation autonomously. The shift to MLLMs in Section III further deepens this integration, allowing UAVs to fuse and interpret streams of visual, sensory, and linguistic data to navigate urban environments, manage swarms, and conduct mission-critical operations such as SAR. Additionally, the technical capabilities and inherent vulnerabilities of LLMs, as outlined in these sections, create specific and urgent ethical imperatives. For example, LLMs are prone to hallucinations, generating plausible yet nonfactual content, which raises concerns about their reliability. This flaw could lead to significant navigation errors or misidentification of targets in the field. Their black-box nature and sophisticated reasoning make it difficult to audit critical UAV decisions. This opacity is compounded in MLLMs, where adversarial attacks can exploit visual vulnerabilities to bypass safety protocols and potentially hijack UAV systems. Moreover, LLMs are trained on large-scale datasets that inevitably contain societal biases; if a model prioritizes disaster relief efforts or the identification of persons of interest, it could systematically disadvantage certain groups. Finally, there are environmental concerns regarding energy consumption and sustainability. Therefore, investigating the ethical implications is critical to the responsible development and deployment of LLM-assisted UAVs. This ensures that the pursuit of technological autonomy is guided by safeguards for safety, fairness, accountability, and societal trust, preventing harm and aligning these systems with human values and safety-critical standards.
V Ethical Challenges and Best Practices for LLM Integration in UAVs
As LLMs and MLLMs become integral to autonomous UAV decision-making, their deployment introduces critical ethical and safety challenges that must be systematically addressed. Privacy risks arise when large-scale datasets include sensitive information collected without explicit consent. Bias and fairness issues may emerge from imbalanced training data, leading to unsafe or discriminatory UAV decisions. Accountability and transparency are further complicated by the black-box nature of LLMs, which hinders failure diagnosis and responsibility attribution. In addition, the high computational cost of LLM training and inference raises sustainability concerns, motivating energy-efficient and environmentally responsible AI practices.

### v2

@@ IV-C MLLM-Assisted UAV Navigation and Collaboration
At a broader level, Liu et al. [68] introduce AerialVLN, a benchmark task that redefines UAV navigation in open, continuous 3D urban spaces. Unlike ground-based VLN, AerialVLN requires UAVs to navigate freely using first-person visual input while handling altitude variation and ambiguous spatial references. This task formalizes key challenges in cross-modal grounding and 3D spatial reasoning that contemporary frameworks such as UAV-VLN, SkyVLN, and AeroDuo aim to address. A comparative summary is provided in Table IX.
Overall, recent advances in VLN demonstrate that tightly integrating LLM-assisted reasoning with visual perception, control-aware execution, and multi-UAV collaboration enables UAVs to translate natural-language instructions into real-time, context-aware flight behavior. These developments establish a strong foundation for future research in scalable, language-driven aerial navigation.
Table IX: Comparison of VLN frameworks for UAVs, highlighting innovations in end-to-end reasoning, safety-aware control, instruction robustness, multi-UAV collaboration, and benchmark development for continuous 3D aerial navigation.
@@
[68] Proposes a new VLN benchmark for UAVs in continuous 3D outdoor urban spaces. Task definition for continuous, open 3D navigation with first-person visual perceptions. Introduces challenges in 3D spatial reasoning, language understanding, and cross-modal alignment. Addresses the gap in aerial-agent navigation for complex urban environments.
@@
IV-D LLM/VLM-Assisted High-Level Mission Planning, Task Reasoning, and Operational Intelligence
@@
Recent research has increasingly focused on enabling UAVs to autonomously interpret visual scenes, understand natural language, and perform time-critical operations. Yaqoot et al. [110] exemplify this trend with UAV-VLRR, a rapid SAR framework that integrates a multimodal vision–language reasoning system using ChatGPT-4o and Molmo with a high-performance NMPC. By directly translating language commands and aerial imagery into actionable target coordinates and executing point-to-point NMPC control with integrated obstacle avoidance, UAV-VLRR eliminates the need for preplanned trajectories. Experimental results demonstrate significant speedups over both human pilots and conventional autopilots, highlighting the effectiveness of combining MLLMs with real-time control for agile, safety-critical missions.
Complementing rapid execution at the control level, Sautenkov et al. [87] extend multimodal reasoning to strategic mission planning with UAV-VLA. Instead of directly controlling low-level flight, their system enables users to specify large-scale aerial missions using natural language, which an LLM interprets as high-level objectives. A vision–language model then identifies relevant targets in satellite imagery, and a final LLM module generates a complete flight plan. Although the resulting trajectories are slightly longer than those designed by expert operators, they are produced more than six times faster, demonstrating the potential of MLLMs to significantly reduce operator workload and streamline global-scale mission design. Beyond navigation and planning, multimodal AI has also demonstrated strong potential for real-time incident detection and situational awareness. Eesaar et al. [24] present an autonomous UAV system that integrates a DJI Matrice 30T with an end-to-end multimodal pipeline for emergency response. In this system, YOLOv12n detects accidents and fires in real-time aerial video, LLaVA-OneVision-Qwen2 generates detailed scene-level descriptions, and GPT-4o mini produces concise natural-language summaries for first responders. Field validation shows that this pipeline can rapidly transform raw aerial imagery into actionable alerts, significantly improving response times and situational awareness in transportation and emergency management scenarios.
@@
IV-E LLM/VLM-Driven Autonomous Control and Real-Time Flight Decision-Making
LLMs and VLMs in UAV systems pose a critical challenge: enabling autonomous agents to accurately infer human intent, interpret complex, dynamic environments, and make safe, context-aware decisions despite the uncertainty and potential hallucinations inherent in modern foundation models.
@@ IV-E LLM/VLM-Driven Autonomous Control and Real-Time Flight Decision-Making
Building on the theme of reliability while emphasizing human collaboration, Krupavs et al. [51] introduce a multimodal architecture for human–machine collaborative UAV navigation. Their system uses GPT-4.1-nano for multimodal reasoning and generates navigation commands, natural-language explanations, and confidence scores, which are presented to operators via a touch-based approval interface. This design enables a closed-loop interaction paradigm in which operators can accept, reject, or modify model-generated actions. Experimental results show high levels of user trust and explainability, with approximately 90% approval of AI-generated instructions. However, observed response latency and imperfect calibration between confidence scores and actual performance highlight ongoing challenges in deploying explainable VLMs for real-time aerial control.
@@
Extending these ideas to goal-directed navigation, Zhang et al. [115] propose VLFly, a unified vision–language navigation framework explicitly designed for UAVs that avoids the common reliance on discrete action spaces and external localization systems. Instead, VLFly directly outputs continuous velocity commands from monocular camera input, integrating an LLM-assisted instruction encoder, a VLM-based goal retriever, and a trajectory planner to generate real-time waypoints. Evaluated across diverse simulation environments without task-specific fine-tuning, VLFly consistently outperforms previous approaches. Real-world flight experiments in both indoor and outdoor settings further confirm its robust open-vocabulary grounding and its ability to interpret abstract or indirect language instructions, highlighting its suitability for zero-shot deployment in unconstrained UAV navigation scenarios. Table X summarizes challenges and solution directions in LLM/VLM-Based UAV Autonomy.
@@
Table X: Key challenges in LLM/VLM-based UAV autonomy—including model uncertainty, interpretability, resource constraints, environmental perception, and intent understanding—paired with solution approaches, required capabilities, and desired operational outcomes.
@@
IV-F Semantic Understanding, Analysis, and Interpretation
As multimodal large models gain traction in UAV-related domains, recent research has begun to explore their potential beyond navigation and control, extending into higher-level reasoning for security and remote sensing applications. Fig. 9 visualizes LLM/VLM-assisted semantic Understanding, Analysis, and Interpretation. Lei et al. [53] exemplify this shift in low-altitude airspace security by proposing a system that integrates MLLMs with heterogeneous perception data to infer the intent of non-cooperative UAVs. By fusing inputs from RGB cameras, infrared sensors, and radar, the system constructs structured representations, which are processed by an MLLM that leverages contextual knowledge and generative reasoning to assess adversarial behavior and recommend defensive actions. A confrontation scenario demonstrates the feasibility of this approach. The authors also identify key challenges – including robust communication, scalability to multi-UAV scenarios, and effective edge–cloud collaboration – that outline future directions for intelligent and adaptive airspace defense architectures.
@@ IV-F Semantic Understanding, Analysis, and Interpretation
IV-G Key Findings and Insights
This section consolidates key observations and empirical findings on the deployment of Multimodal Large Language Models (MLLMs) in UAV applications. Emphasis is placed on quantitative evidence, benchmark results, and UAV-specific limitations that directly affect real-world reliability, efficiency, and safety as a result.
IV-G1 Benchmarking MLLMs for Multi-UAV Collaborative Perception
AirCopBench [113] is the first benchmark explicitly designed to evaluate MLLMs in multi-UAV collaborative perception under degraded sensing conditions. The benchmark comprises over 2,900 multi-view aerial images and more than 14,600 VQA samples, spanning four task categories (Scene Understanding, Object Understanding, Perception Assessment, and Collaborative Decision) and 14 task types. Evaluation of 40 state-of-the-art MLLMs reveals a substantial performance gap: the best-performing model achieves only 59.17% accuracy, lagging human performance by more than 24% on average, with large variance across task categories. Fine-tuning results indicate that sim-to-real transfer can partially mitigate these gaps, highlighting the importance of UAV-specific data and training strategies for collaborative aerial perception.
IV-G2 Speculative Decoding for Efficient LLM Inference
Inference latency remains a critical bottleneck for UAV deployment, as conventional autoregressive decoding requires one forward pass per generated token. Speculative decoding improves efficiency by allowing multiple candidate tokens to be generated and verified in parallel within a single forward pass, significantly reducing end-to-end latency while preserving output correctness. Recent approaches, such as EAGLE-3, further eliminate the need for a separate draft model by leveraging the target model’s internal hidden states. Although primarily evaluated in general-purpose LLM settings, these techniques are highly relevant to UAV platforms, where real-time responsiveness and energy efficiency are tightly constrained.
IV-G3 Multi-Agent Framework for MLLMs Training
@@
IV-G4 Balancing Efficiency and Reasoning Diversity in Vision-Language Models for UAV Applications
@@
Results in [35] comprising 50,019 samples organized into progressively challenging stages, further highlight that domain-specific data can outweigh sheer model scale, outperforming the posed UAV-VL-R1 model, which achieves strong reasoning performance while maintaining a compact footprint suitable for onboard deployment. Ablation studies reveal a key trade-off: while supervised fine-tuning (SFT) accelerates convergence and stabilizes early performance, excessive reliance on SFT reduces reasoning diversity, particularly in counting and numerical inference tasks. In contrast, the multi-stage GRPO framework enhances reasoning diversity and cross-task generalization by comparing intra-group rewards. The associated HRVQA-VL dataset, consisting of 50,019 samples organized into progressively challenging stages, further highlights that domain-specific data can outweigh sheer model scale, outperforming both 2B and 72B Qwen2-VL baselines under UAV constraints.
IV-G5 Understanding and Evaluating Hallucination Behavior in Aerial LVLMs
@@
Basak et al. [10] introduce AeroCaps, the first aerial-view image captioning dataset, comprising 1,256 human-annotated images. Their analysis shows that aerial perspectives significantly increase hallucination rates, especially for object identity, spatial relations, and numerical descriptions, with error frequency increasing with caption length. Notably, mitigation strategies such as object hints and LLM-based self-evaluation are only partially effective, underscoring the difficulty of reliable hallucination detection. These findings indicate that current LVLMs are insufficiently robust for real-world UAV deployment without further advances in dataset scale, rigorous evaluation, and grounding-aware model design. The study emphasizes that, despite impressive performance on conventional benchmarks, current LVLMs are not yet sufficiently robust or trustworthy for deployment in real-world aerial applications, including automated surveillance, environmental monitoring, and disaster response. The results suggest that progress in this area will require not only improved model architectures and training strategies but also domain-specific datasets, rigorous evaluation protocols, and transparent reporting of failure modes. As aerial data becomes increasingly central to AI-driven decision-making, addressing these challenges will be critical to ensuring the safe deployment of AI.
V Ethical Challenges and Future Directions